pandasurvive:

Solve for x. Use the quadratic formula. 3x^2+7x+3=0 Enter the solutions, in simplified radical form, in the boxes.

Nnesha:

use the quadratic formula \[\large\rm \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] where a is the leading coefficient b is the coefficient of middle term c =constant \[Ax^2+Bx+C=0\]

pandasurvive:

do i turn 2.70325740 into 2.703 or 2.7?

Nnesha:

all of them are not in radical form

pandasurvive:

how do I put them in radical form?

Nnesha:

plugin the a ,b ,c values into the formula and then simplify it that will give you radical form answer.

pandasurvive:

So a=2 b=1 and c=3?

pandasurvive:

@Nnesha

Nnesha:

c is correct. a is the leading coefficient not the exponent \[Ax^2\] the number infront of the `x^2` variable will be a

Nnesha:

so for your equation A=3 and b=7 \[\color{Red}{3}x^2+\color{blue}{7}x+\color{orange}{3}=0\] \[\color{Red}{A}x^2+\color{blue}{B}x+\color{orange}{C}=0\]

pandasurvive:

so -7+- 49-36 over 8?

Nnesha:

yes that's correct. don't forget the denominator 2(a)=2(3)

pandasurvive:

isnt 2^3= to 8?

pandasurvive:

oh wait its 2*3 lol 6

Nnesha:

\[\frac{ -7 \pm \sqrt{13 }}{ 6 }\]

Nnesha:

yes right and since it's plus minus we can separate tem \[\frac{ -7-\sqrt{13} }{ 6 } , \frac{-7+\sqrt{13}}{6}\]

pandasurvive:

thats simplified radical form?

Nnesha:

that's it.

pandasurvive:

thank you :D

Nnesha:

yw